Xanadu Mining is considering three mutually exclusive alternatives, as shown in the table below. MARR is 10.5%/yea

EOY A001 B002 C003
0 -200 -110 -154
1 95 60 78
2 105 60 78
3 115 60 78
4 125 70 78
Answer #1

ANSWER:

I = 10.5%

Pw of a001 = cash flow in year 0 + cash flow in year 1(p/a,i,n) + increase in cash flow(p/g,i,n)

pw = -200 + 95(p/a,10.5%,4) + 10(p/g,10.5%,4)

pw = -200 + 95 * 3.14 + 10 * 4.31

pw = -200 + 297.91 + 43.14

pw = 141.04

Pw of a002 = cash flow in year 0 + cash from year 1 to year 3(p/a,i,n) + cash flow in year 4(p/f,i,n)

pw = -110 + 60(p/a,10.5%,3) + 70(p/f,10.5%,4)

pw = -110 + 60 * 2.47 + 70 * 0.67

pw = -110 + 147.91 + 46.95

pw = 84.86

Pw of a003 = cash flow in year 0 + cash flow from year 1 to year 4(p/a,i,n)

pw = -154 + 78(p/a,10.5%,4)

pw = -154 + 78 * 3.14

pw = -154 + 244.60

pw = 90.60

since the pw of a001 is the best , therefore it is the best alternative.